Can you help me simplify this problem. (x/3)^3
Found 2 solutions by stanbon, MathLover1:
Answer by stanbon(75887)

simplify this problem. (x/3)^3
----
(x/3)^3 = (x/3)(x/3)(x/3) = x^3/27
-----------------
Cheers,
Stan H.
-------------

Answer by MathLover1(20849)